OpenStudy (anonymous):

PLZ HELP WILL BE REWARDED!!! Which literal equations are equivalent to w = Fd? Choose all answers that are correct. A. B. C. D.

OpenStudy (anonymous):

I will put the options

OpenStudy (anonymous):

ok idk this math im sorry

OpenStudy (anonymous):

A. Fw B. w/F C.F=w/d D. F=d/w

OpenStudy (anonymous):

i will hjelp

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

its a multiple choice questions so most likely there's 2 answers.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wel, i don't think there are moer

OpenStudy (anonymous):

ok thx and are you sure tht the answer is c? Just to make sure lol

OpenStudy (anonymous):

@StudyGurl14 may you help?

OpenStudy (studygurl14):

C is the only one that looks correct, but it looks like there is missing information for answer choices A and B...

OpenStudy (studygurl14):

With these type of questions, there are almost always at least two correct answers...so are you sure you typed A and B correctly @Cholo123 ?

OpenStudy (anonymous):

hold on plz i will check

OpenStudy (anonymous):

yes there are sorry A. d=Fw B. d=w/F C. F=w/d D. F=d/w

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@ttb123456789

OpenStudy (anonymous):

@StudyGurl14 am i right if i say tht the answers r a and b?

OpenStudy (studygurl14):

No. the answers are b and c

OpenStudy (anonymous):

oh ok thx do u mind if i ask one more question?

OpenStudy (studygurl14):

Sure, and medal please. (only one more in this thread. more than that creates long threads, which cause lagging)

OpenStudy (anonymous):

ok well i actually have 2 question one i will post here and one i will tage u in the new thread Which literal equation is equivalent to d = 0.2t? A. B. C. D.

OpenStudy (anonymous):

ugh i will message the options

OpenStudy (anonymous):

A. t=d/0.2 B. t=0.2/d C. 0.2=t/d D. 0.2=dt

OpenStudy (anonymous):

@StudyGurl14 sorry it took forever lol

OpenStudy (studygurl14):

\(\large d=0.2t\rightarrow t=\frac{d}{0.2}\)

OpenStudy (anonymous):

so A?

OpenStudy (studygurl14):

yea

OpenStudy (anonymous):

or B?

OpenStudy (anonymous):

ok thx and i will tag u in the new thread

OpenStudy (studygurl14):

ok.